# On the heat kernel of non-minimal second-order operators

### v1

@@ V Final result
First of all, by using Eq.s (19), (10), (15) and (52) we can write down the divergent part due to the box operator as
@@ V.1 On our computational tools
Due to the generality and complicated nature of our calculation, the final expression of the divergent part of Eq. (25) is still very long. Thus, in order to avoid typos in the transcription, we have chosen to directly print out the resulting scalar contraction from the code using TexPrint. This implies that the vector bundle indices are displayed explicitly as
@@
where $\phi_{\mu\nu}$ is the quantum field that whose fluctuations we are taking into account. Then, we have that
@@ V.2 First part of the result
where the numerical subscripts indicate the mass-dimension of the tensors. Moreover, mass-dimension four terms are always total-derivative endomorphisms, thence we discard them. Finally, $\hat{M}_{3}$ can only enter the result when considering terms linear in the derivative interaction tensor $\hat{V}$.
@@ V.3.1 Order $\hat{V}^{0}$ terms
Thus, if one is interested in deriving the back-reaction of some matter fields on the pseudo-Riemannian structure, then, if the assumption of Subsect. III.1 are met, it is sufficient to take into account this last equation, together with Eq.s (V.2), (67) and (V.2).
@@ V.3.5 Order $\hat{V}^{4}$ terms
Finally, terms that are quartic in $\hat{V}$ can be written in the following simple fashion
@@ VI.1 Vector field
Let us consider a Lagrangian for a vector field $A_{\mu}$ which comprises a non-minimal kinetic term, as well as cubic and quartic-self interactions. Specifically, we parameterize it as
@@
For example, we may think of this Lagrangian to describe the torsion vector Sauro:2022chz ; Sauro:2022hoh ; Paci:2024ohq or Weyl vectors Ghilencea:2018dqd in a metric-affine geometry. Alternatively, such kind of dynamics may also arise by considering vector fields coupled to scalars as in Buchbinder:2024fwy . The resulting second-order differential operator which acts on contra-variant vectors $A^{\nu}$ takes the following form
$F^{\mu}{}_{\nu}(\nabla)=\square\delta^{\mu}{}_{\nu}-\lambda\nabla^{\mu}\nabla_{\nu}-\tfrac{f}{3}\left(A_{\nu}\nabla^{\mu}-A^{\mu}\nabla_{\nu}\right)-bR^{\mu}{}_{\nu}-cR\delta^{\mu}{}_{\nu}-gA^{\mu}A_{\nu}\,.$ (76)
@@
In analogy with Eq. (76), we fix the ambiguity in ordering the covariant derivatives of $\hat{K}$ by putting that which acts on $A^{\nu}$ to the right. The auxiliary integration parameter $\zeta$ is introduced by simply letting $\lambda\rightarrow\lambda\zeta$. Then, the final form of the trace of the second Seeley-DeWitt coefficient is found by inserting Eq.s (77) into Eq.s (67), (V.2), (V.2), (69), (71), (70), (123), (72), (C.2), (124), (C.2), (73) and (74) and computing the $\zeta$ integrals. Furthermore, we parameterize the result in terms of $\gamma=\frac{\lambda}{1+\lambda}$. Accordingly, we have that
@@ VI.2 Kalb-Ramond field
As in the previous example, we have decided to sort the covariant derivatives such that the rightmost one generates divergences of $B^{\rho\sigma}$. Both $\hat{K}$ and $\hat{M}_{2}$ are singular for $b\zeta=-2$, which is exactly the value for which the Lagrangian (79) propagates only the transverse modes. Due to the linear dependence on $\zeta$ in $D^{\alpha\beta\,\,\mu\nu}{}_{\rho\sigma}$ this example falls into the simplified category of non-minimal operators dealt with in Subsect. III.1. Thus, to find contribution to the trace of $\hat{a}_{2}$ due to the non-minimal part of the operator we have to insert the tensors written in the previous equations (81) into Eq.s (67), (V.2) and (V.2). Performing the $\zeta$ integral we find
@@
Next, we consider the remaining part of the trace-log of the operator (80), i.e., that which depends on the endomorphism. By using Eq. (69) we obtain the contribution to the trace of the second Seeley-DeWitt coefficient, and the complete form of the trace of this coefficient is given by
@@ VI.3 Torsion field: a simple toy model
While a systematic derivation of the phase diagram in the most general metric-torsion theory would settle a common agreement regarding the physical properties of these theories, this is not the aim of the present discussion. Indeed, our aim is to provide a simple example for showing how to treat the presence of non-minimal kinetic terms in torsion theories, and to point out some phenomenological consequences of taking them into account. To this end, we consider the following Lagrangian
@@
Once again, we have sorted the covariant derivatives to act with divergences on $T^{\rho}{}_{\mu\nu}$. Clearly, the last equation become singular as $c\,\zeta\rightarrow-2$. Next, we promote momentum space-expressions to coordinate-based ones, and we compute the $\hat{M}_{2}$ differential operator, whose tensor coefficients are given by the following equation
@@
Now we have all the necessary ingredients to compute the trace of the second Seeley-DeWitt coefficient of the non-minimal operator Eq. (89). First, we focus on the trace-log of the non-minimal part of the operator (89). To this end, we insert Eq.s (90) and (91) in the general result, which given by the sum of Eq.s  (67), (V.2) and (V.2). By performing the $\zeta$ integral we obtain the following contribution to the trace of $\hat{a}_{2}$
@@
Therefore, in strike contrast with the examples of electrodynamics Barvinsky:1985an and of the Kalb-Ramond field the dependence on the non-minimal parameter $c$ does not disappear. Accordingly, we cannot treat the Lagrangian Eq. (87) as describing a gauge theory of a vector-valued $2$-form as a consistent quantum field theory on a curved background. Consequently, taking into account the quantum fluctuations of the metric field becomes mandatory in order to understand if this theory generates a quantum anomaly. However, due to the first Bianchi identities, we cannot couple the generalized field strength ${\cal{F}}^{\rho}{}_{\lambda\mu\nu}$ to the Riemann tensor, which yields a significantly different behavior w.r.t. that of the vector theory treated in Subsect. VI.1.
@@ VII Discussion and outlooks
In the present paper we have derived for the first time a general expression for the local part of the trace of the second Seeley-DeWitt coefficient of a non-minimal operator in $d=4$. To this end we have employed the trick of Barvinsky:1985an for re-writing the trace-log of a non-minimal operator as the integral over an auxiliary parameter of a sum of functional traces. However, we have followed Melichev:2025hcg in splitting the calculation into two different parts, in such a way that only the first one involves the integration over the auxiliary parameter. As we have discussed at the end of Sect. III such a split significantly improves the computational speed for two reasons. On the one hand, the integrals that have to be computed usually are way less that in the method of Barvinsky:1985an ; on the other hand, we generate symmetrized metrics $g^{(n)}$ only up to $n=6$, compared to $n=8$ in Barvinsky:1985an ’s method. Thus, we have obtained a reduction of a factor $\tfrac{1}{13\times 15}$ in the number of generated tensor structure at order $\hat{V}^{4}$, and analogous improvements appear at all orders in $\hat{V}$. All of the model-independent formulae for the local part of the trace of the second Seeley-DeWitt coefficient are also attached to this paper in a Mathematica notebook.
@@
In Sect. VI we have applied the general result to three simple cases. Our findings are compatible with known results in the literature Barvinsky:1985an ; Melichev:2025hcg ; Sangy:2025jxk . An interesting future development of these applications is to derive the divergent part of the effective action for higher-spins, torsion and non-metricity fields. This would provide a comprehensive picture of the radiative effects of these quantum fields on the high-energy behavior of higher-derivative gravity. In particular, it would be possible to analyze which parts of the physical parameter space are compatible with asymptotic safety (see, e.g., Melichev:2025hcg for a recent effort in this direction). Part of this program was carried out in Piva:2021nyj , though the author did not consider non-minimal terms for higher-spins, while the radiative (in-)stability of torsion and non-metricty theories was studied in Marzo:2021iok .
As we have mentioned in the Introduction I, there exists another way of deriving the trace-log of non-minimal operators within the heat kernel method, which relies on spin-parity decompositions Groh:2011dw . Such a program was carried out on a maximally symmetric background in Martini:2023apm integrating out torsion fluctuations, employing the spin-parity decomposition developed in Martini:2023rnv . In general, if the (bosonic) quantum field that we are integrating out has spin $s$, we will be required to know the precise form of the traced heat kernel coefficients of minimal operators up to order 2+2s2+2\,s. Therefore, while pursuing this route might be preferable when dealing with the torsion or non-metricity tensors, we argue that it will be less convenient for fields of spin $s\geq 4$.
@@ VIII Acknowledgements
The author whishes to thank Oleg Melichev for useful comments and interesting discussions related to the topic of the present paper.
@@ Appendix A Commutator identities
Let us exploit a Laplace transform to write down the inverse of the n-th power of $\hat{\square}$, i.e.,
@@
By using this integral representation we can trade the commutation of $\hat{\square}^{-1}$ with the commutation of $\hat{\square}$, thus keeping under control the mass-dimension of the operators that we are dealing with. Indeed, the following set of equations hold
@@
where the ellipsis indicate that we are discarding higher-dimensional terms. In doing so we are tacitly assuming that $\hat{M}$ has at dimension one, whence triple commutators would at most give rise to boundary terms.
@@ Appendix B Coincidence limits
In this section we present all the coincidence limits up to dimension four operators of the Synge world function, the reduced Van-Vleck determinant and the Seeley-DeWitt coefficients. As we have done in the main text, we discard surface terms. The Seeley-DeWitt coefficients are those of the minimal operator
@@ B.2 Covariant derivatives of the VanVleck-Morette determinant
Finally, for dimension four terms we make use of computer algebra, and we obtain
@@ B.3 Covariant derivatives of the Seeley-DeWitt coefficient
After having derived the coincidence limits of the Synge world function and VanVleck-Morette determinant up to dimension four terms, we focus on the Seeley-DeWitt coefficients. The zero-th coefficient $\hat{a}_{0}$ is interpreted as a parallel displacement operator Barvinsky:1985an . As such, its covariant Taylor expansion only depends on the pseudo-Riemannian structure and on the generalized curvature $\hat{\mathcal{R}}_{\mu\nu}$, and it cannot depend on the specific differential operator that we are dealing with. This geometrical interpretation and the initial conditions of the Cauchy system Eq. (8) yield
@@
By applying one more derivative and using Eq.s (103) and (104) we find the first non-trivial result

### v2

@@ V Final result
First of all, by using Eqs. (19), (10), (15) and (52) we can write down the divergent part due to the box operator as
@@ V.1 On our computational tools
Due to the generality and complicated nature of our calculation, the final expression of the divergent part of Eq. (25) is still very long. Thus, in order to avoid typos in the transcription, we have chosen to directly print out the resulting scalar contraction from the code using TexPrint, and to attach the code itself to the arXiv version of the paper Sauro:2025nb . This implies that the vector bundle indices are displayed explicitly as
@@
where $\phi_{\mu\nu}$ is the quantum field whose fluctuations we are taking into account. Then, we have that
@@ V.2 First part of the result
where the numerical subscripts indicate the mass-dimension of the tensors. Moreover, mass-dimension-four terms are always total-derivative endomorphisms, thence we discard them. Finally, $\hat{M}_{3}$ can only enter the result when considering terms linear in the derivative interaction tensor $\hat{V}$.
@@ V.3.1 Order $\hat{V}^{0}$ terms
Thus, if one is interested in deriving the back-reaction of some matter fields on the pseudo-Riemannian structure, then, if the assumption of Subsect. III.1 are met, it is sufficient to take into account this last equation, together with Eqs. (V.2), (67) and (V.2).
@@ V.3.5 Order $\hat{V}^{4}$ terms
Finally, terms that are quartic in $\hat{V}$ can be written in the simple fashion
@@ V.4 Limiting case
V.5 Some words on other computational schemes
Recently, a new computational approach towards non-minimal differential operators was proposed in Barvinsky:2025jbw , whose main concern is to derive the low-energy part of the heat kernel in an exact way. It relies on the introduction of spin-projectors, which are employed in a different manner with respect to previous studies Martini:2023apm . Indeed, in flat space these projectors allow to express the untraced heat kernel as a sum over the spin-parity eigenspaces. When the background geometry is curved the projectors are replaced by quasi-projectors, and an appropriate subtraction procedure is introduced to deal with the non-locality of them. A systematic perturbative expansion in the background dimension is then carried out on the full heat kernel, which is free of IR pathologies.
When some simple operators are considered and a judicious choice of the endomorphism is made the approach of Barvinsky:2025jbw yields an exact expression for the full untraced heat kernel. However, for more complicated differential operators this technique requires to deal with many integrations over proper time of some complicated nested commutators. The final expressions shall still be non-perturbative in the proper time, but they do not allow to implement the final result on a computer to yield an easy-to-use expression. Nevertheless, this result represents a cornerstone in the derivation of infrared properties of $1$-loop effective action stemming from non-minimal second order operators.
On the other hand, the approach that we have followed in this paper was to focus on the traced heat kernel coefficients, and in particular on the second one. By doing this and employing the results of Sect. IV on universal functional traces we were able to provide a model-independent expression for this coefficient. Furthermore, the open-source implementation in Mathematica of this Seeley-DeWitt coefficient significantly contributes to the applicability of our results. Finally, no subtraction procedure is needed in this case, since we are only interested in the UV divergent part of the effective action. Therefore, this paper and Barvinsky:2025jbw provide two complementary results on the heat kernel of non-minimal second-order operators.
@@ VI.1 Vector field
Let us consider a Lagrangian for a vector field $A_{\mu}$ which comprises a non-minimal kinetic term, as well as cubic and quartic self-interactions. Specifically, we parameterize it as
@@
For example, we may think of this Lagrangian to describe the torsion vector Sauro:2022chz ; Sauro:2022hoh ; Paci:2024ohq or Weyl vectors Ghilencea:2018dqd in a metric-affine geometry. Alternatively, such kind of dynamics may also arise by considering vector fields coupled to scalars as in Buchbinder:2024fwy . The beta functions of higher-derivative gravity due to a non-minimal vector field’s quantum fluctuations were also computed in Piva:2023noi : the author did not consider the cubic interaction, but mixed loops of gravitons and vectors were taken into account. The resulting second-order differential operator which acts on contra-variant vectors $A^{\nu}$ takes the form
$F^{\mu}{}_{\nu}(\nabla)=\square\delta^{\mu}{}_{\nu}-\lambda\nabla^{\mu}\nabla_{\nu}-\tfrac{f}{3}\left(A_{\nu}\nabla^{\mu}-A^{\mu}\nabla_{\nu}\right)-bR^{\mu}{}_{\nu}-cR\delta^{\mu}{}_{\nu}-gA^{\alpha}A_{\alpha}\delta^{\mu}{}_{\nu}-2gA^{\mu}A_{\nu}\,.$ (76)
@@
In analogy with Eq. (76), we fix the ambiguity in ordering the covariant derivatives of $\hat{K}$ by putting that which acts on $A^{\nu}$ to the right. The auxiliary integration parameter $\zeta$ is introduced by simply letting $\lambda\rightarrow\lambda\zeta$. Then, the final form of the trace of the second Seeley-DeWitt coefficient is found by inserting Eqs. (77) into Eq.s (67), (V.2), (V.2), (69), (71), (70), (123), (72), (C.2), (124), (C.2), (73) and (74) and computing the $\zeta$ integrals. Furthermore, we parameterize the result in terms of $\gamma=\frac{\lambda}{1+\lambda}$. Accordingly, we have that
@@ VI.2 Kalb-Ramond field
As in the previous example, we have decided to sort the covariant derivatives such that the rightmost one generates divergences of $B^{\rho\sigma}$. Both $\hat{K}$ and $\hat{M}_{2}$ are singular for $b\zeta=-2$, which is exactly the value for which the Lagrangian (79) propagates only the transverse modes. Due to the linear dependence on $\zeta$ in $D^{\alpha\beta\,\,\mu\nu}{}_{\rho\sigma}$ this example falls into the simplified category of non-minimal operators dealt with in Subsect. III.1. Thus, to find contribution to the trace of $\hat{a}_{2}$ due to the non-minimal part of the operator we have to insert the tensors written in the previous equations (81) into Eqs. (67), (V.2) and (V.2). Performing the $\zeta$ integral we find
@@
Next, we consider the remaining part of the trace log of the operator (80), i.e., that which depends on the endomorphism. By using Eq. (69) we obtain the contribution to the trace of the second Seeley-DeWitt coefficient, and the complete form of the trace of this coefficient is given by
@@ VI.3 Torsion field: A simple toy model
While a systematic derivation of the phase diagram in the most general metric-torsion theory would settle a common agreement regarding the physical properties of these theories, this is not the aim of the present discussion. Indeed, our aim is to provide a simple example for showing how to treat the presence of non-minimal kinetic terms in torsion theories, and to point out some phenomenological consequences of taking them into account. To this end, we consider the Lagrangian
@@
Once again, we have sorted the covariant derivatives to act with divergences on $T^{\rho}{}_{\mu\nu}$. Clearly, the last equation becomes singular as $c\,\zeta\rightarrow-2$. Next, we promote momentum-space expressions to coordinate-based ones, and we compute the $\hat{M}_{2}$ differential operator, whose tensor coefficients are given by the following equation
@@
Now we have all the necessary ingredients to compute the trace of the second Seeley-DeWitt coefficient of the non-minimal operator Eq. (89). First, we focus on the trace log of the non-minimal part of the operator (89). To this end, we insert Eqs. (90) and (91) in the general result, which given by the sum of Eqs. (67), (V.2) and (V.2). By performing the $\zeta$ integral we obtain the following contribution to the trace of $\hat{a}_{2}$
@@
Therefore, in strike contrast with the examples of electrodynamics Barvinsky:1985an and of the Kalb-Ramond field the dependence on the non-minimal parameter $c$ does not disappear. Accordingly, we cannot treat the Lagrangian Eq. (87) as describing a gauge theory of a vector-valued $2$-form as a consistent quantum field theory on a curved background. Consequently, taking into account the quantum fluctuations of the metric field becomes mandatory in order to understand if this theory generates a quantum anomaly. However, due to the first Bianchi identities, we cannot couple the generalized field strength ${\cal{F}}^{\rho}{}_{\lambda\mu\nu}$ to the Riemann tensor, which yields a significantly different behavior with respect to that of the vector theory treated in Subsect. VI.1.
@@ VII Discussion and outlooks
In this paper we have derived for the first time a general expression for the local part of the trace of the second Seeley-DeWitt coefficient of a non-minimal operator in $d=4$. To this end we have employed the trick of Barvinsky:1985an for rewriting the trace log of a non-minimal operator as the integral over an auxiliary parameter of a sum of functional traces. However, we have followed Melichev:2025hcg in splitting the calculation into two different parts, in such a way that only the first one involves the integration over the auxiliary parameter. As we have discussed at the end of Sect. III such a split significantly improves the computational speed for two reasons. On the one hand, the integrals that have to be computed usually are way less that in the method of Barvinsky:1985an ; on the other hand, we generate symmetrized metrics $g^{(n)}$ only up to $n=6$, compared to $n=8$ in Barvinsky:1985an ’s method. Thus, we have obtained a reduction of a factor $\tfrac{1}{13\times 15}$ in the number of generated tensor structure at order $\hat{V}^{4}$, and analogous improvements appear at all orders in $\hat{V}$. All of the model-independent formulae for the local part of the trace of the second Seeley-DeWitt coefficient are also attached to this paper in a Mathematica notebook Sauro:2025nb .
@@
In Sect. VI we have applied the general result to three simple cases. Our findings are compatible with known results in the literature Barvinsky:1985an ; Melichev:2025hcg ; Sangy:2025jxk . An interesting future development of these applications is to derive the divergent part of the effective action for higher-spins, torsion and non-metricity fields. This would provide a comprehensive picture of the radiative effects of these quantum fields on the high-energy behavior of higher-derivative gravity. In particular, it would be possible to analyze which parts of the physical parameter space are compatible with asymptotic safety (see, e.g., Melichev:2025hcg for a recent effort in this direction). Part of this program was carried out in Piva:2021nyj , though the author did not consider non-minimal terms for higher-spins, while the radiative (in)stability of torsion and non-metricty theories was studied in Marzo:2021iok .
As we have mentioned in the Introduction I, there exists another way of deriving the trace log of non-minimal operators within the heat kernel method, which relies on spin-parity decompositions Groh:2011dw . Such a program was carried out on a maximally symmetric background in Martini:2023apm integrating out torsion fluctuations, employing the spin-parity decomposition developed in Martini:2023rnv . In general, if the (bosonic) quantum field that we are integrating out has spin $s$, we will be required to know the precise form of the traced heat kernel coefficients of minimal operators up to order 2+2s2+2\,s. Therefore, while pursuing this route might be preferable when dealing with the torsion or non-metricity tensors, we argue that it will be less convenient for fields of spin $s\geq 4$.
@@ VIII Acknowledgements
The author whishes to thank Oleg Melichev for useful comments and interesting discussions related to the topic of this paper.
@@ Appendix A Commutator identities
Let us exploit a Laplace transform to write down the inverse of the $n$th power of $\hat{\square}$, i.e.,
@@
By using this integral representation we can trade the commutation of $\hat{\square}^{-1}$ with the commutation of $\hat{\square}$, thus keeping under control the mass-dimension of the operators that we are dealing with. Indeed, the following set of equations holds
@@
where the ellipsis indicates that we are discarding higher-dimensional terms. In doing so we are tacitly assuming that $\hat{M}$ has dimension one, whence triple commutators would at most give rise to boundary terms.
@@ Appendix B Coincidence limits
In this section we present all the coincidence limits up to dimension-four operators of the Synge world function, the reduced Van-Vleck determinant and the Seeley-DeWitt coefficients. As we have done in the main text, we discard surface terms. The Seeley-DeWitt coefficients are those of the minimal operator
@@ B.2 Covariant derivatives of the VanVleck-Morette determinant
Finally, for dimension-four terms we make use of computer algebra, and we obtain
@@ B.3 Covariant derivatives of the Seeley-DeWitt coefficient
After having derived the coincidence limits of the Synge world function and VanVleck-Morette determinant up to dimension-four terms, we focus on the Seeley-DeWitt coefficients. The zeroth coefficient $\hat{a}_{0}$ is interpreted as a parallel displacement operator Barvinsky:1985an . As such, its covariant Taylor expansion only depends on the pseudo-Riemannian structure and on the generalized curvature $\hat{\mathcal{R}}_{\mu\nu}$, and it cannot depend on the specific differential operator that we are dealing with. This geometrical interpretation and the initial conditions of the Cauchy system Eq. (8) yield
@@
By applying one more derivative and using Eqs. (103) and (104) we find the first non-trivial result